In a fluorescent tube of diameter 3.6 cm , 1.1 × 1018 electrons (with a charge of −e) and 2.3 × 1018 positive ions (with a charge of +e) flow in opposite directions through a cross-sectional each second. What is the current in the tube? The fundamental charge is 1.602 × 10−19 C. Answer in units of A.
Answer #1

Net current flowing through tube is due to elections and due to flow of positive ions . current due to electrons flow opposite to the direction of flow of electrons, hence net current is in the direction of positive ions.

Inet= Ie+ Iions =(1.6×10-19 ×1.1×1018)C/s+(1.6×10-19×2.3×1018)C/s=0.176A+0.368A=0.544 A. Answer.